# PowerGraph: A power grid benchmark dataset for graph neural networks

### v2

@@ A.4 Class targeted explanations
For benchmarking explanations in section 4.1, we focus on explaining Category A graphs of the multi-class problem, i.e., the power grids that fail to serve the demand (DNS>0). The objective is to shed light on the tripped lines after the first contingency. We use the multi-class problem rather than the binary classification problem that classifies states according to the demand not served (DNS) only, i.e. distinguishes power grids that serve the demand (DNS=0, label 1) from the ones that do not (DNS>0, label 0). In the multi-class problem, the model learns to distinguish cascading failure scenarios, while in the binary setting, Category A and B are considered the same type of grids (class DNS>0). Explaining DNS>0 in the multi-class problem allows us to focus on the case where some lines are tripped when DNS>0 and, therefore, expect the model to learn the cascading edges for this class of grids.
A.5 Balanced accuracy
@@ C.1 Dataset documentation and intended uses
  • edge_attr.mat: edge feature matrix for the power flow problem
  • edge_attr_opf.mat: edge feature matrix for the optimal power flow problem
  • edge_index.mat: branch list for the power flow problem
  • edge_index_opf.mat: branch list for the optimal power flow problem
  • X.mat: node feature matrix for the power flow problem
  • Xopf.mat:node feature matrix for the optimal power flow problem
  • Y_polar.mat: node output matrix for the power flow problem
  • Y_polar_opf.mat: node output matrix for the optimal power flow problem
@@
  • blist.mat: branch list also called edge order or edge index
  • of_bi.mat: binary classification
  • of_reg.mat: regression labels
  • of_mc.mat: multi-class labels
  • Bf.mat: node feature matrix
  • Ef.mat: edge feature matrix
  • exp.mat: ground-truth explanation

### v3

@@ A.4 Class targeted explanations
For benchmarking explanations in section 4, we focus on explaining Category A graphs of the multi-class problem, i.e., the power grids that fail to serve the demand (DNS>0). The objective is to shed light on the tripped lines after the first contingency. We use the multi-class problem rather than the binary classification problem that classifies states according to the demand not served (DNS) only, i.e. distinguishes power grids that serve the demand (DNS=0, label 1) from the ones that do not (DNS>0, label 0). In the multi-class problem, the model learns to distinguish cascading failure scenarios, while in the binary setting, Category A and B are considered the same type of grids (class DNS>0). Explaining DNS>0 in the multi-class problem allows us to focus on the case where some lines are tripped when DNS>0 and, therefore, expect the model to learn the cascading edges for this class of grids.
A.5 Reconstructing temporal graphs for in-depth analysis of cascading failures in PowerGraph
Each instance of the graph-level dataset in PowerGraph, designed for cascading failure analysis, is accompanied by an explainability mask that encodes the stages of the cascading failure process. This mask is stored in a MATLAB structure exp.mat, see in Appendix C.1. Each entry in the MATLAB structure is a vector containing all failures at all stages, with the order of the vector corresponding to the sequence of line failures in the cascading event. Using this information, a temporal graph can be reconstructed from PowerGraph.
A.6 Balanced accuracy
@@ C.1 Dataset documentation and intended uses
  • edge_attr.mat: edge feature matrix for the power flow problem (branch conductance $G_{ij}$, branch susceptance $B_{ij}$.)
  • edge_attr_opf.mat: edge feature matrix for the optimal power flow problem (branch conductance $G_{ij}$, branch susceptance $B_{ij}$.)
  • edge_index.mat: list of branches, represented as connection from node - to node.
  • edge_index_opf.mat: list of branches, represented as connection from node - to node.
  • X.mat: node feature matrix for the power flow problem (active power generation $P_{g}$ active power demand $P_{d}$, reactive power generation $Q_{g}$ reactive power demand $Q_{d}$, voltage magnitude $V$ , and voltage angle $\theta$, the number of loads $N_{loads}$, and number of generators $N_{gen}$).
  • Xopf.mat:node feature matrix for the optimal power flow problem (active power generation $P_{g}$ active power demand $P_{d}$, reactive power generation $Q_{g}$ reactive power demand $Q_{d}$, voltage magnitude $V$ , and voltage angle $\theta$, the number of loads $N_{loads}$, and number of generators $N_{gen}$).
  • Y_polar.mat: node output matrix for the power flow problem (active power generation $P_{g}$ active power demand $P_{d}$, reactive power generation $Q_{g}$ reactive power demand $Q_{d}$, voltage magnitude $V$ , and voltage angle $\theta$).
  • Y_polar_opf.mat: node output matrix for the optimal power flow problem (active power generation $P_{g}$ active power demand $P_{d}$, reactive power generation $Q_{g}$ reactive power demand $Q_{d}$, voltage magnitude $V$ , and voltage angle $\theta$).
@@
  • blist.mat: list of branches, represented as connection from node - to node
  • of_bi.mat: binary classification labels ($DNS=0$ or $DNS\neq 0$)
  • of_reg.mat: regression labels ($DNS$)
  • of_mc.mat: multi-class classification labels (See Table 2)
  • Bf.mat: node feature matrix (Net active power at bus $P_{net}$, Net apparent power at bus $S_{net}$, Voltage magnitude $V$
  • Ef.mat: edge feature matrix (Active power flow $P_{i,j}$, Reactive power flow $Q_{i,j}$, Line reactance $X_{i,j}$, Line rating $lr_{i,j}$)
  • exp.mat: ground-truth explanation (See A.5)